8 = 2^3
I submitted the first problem incorrectly and tried to solve the problem with some understanding from the text. I states if x =a^p then 1/x =1/ap = a-p. So I was trying to decipher the above example using the formula.
8= 2^3 then 1/2 = 0.5 = 8- 1/3. I wanted to know how the 1/2 was arrived at.Or can someone explain the process. I am Clueless
Thanking you in advance
TeiTei

Answer by kev82(148)

If x=a^p then 1/x=1/(a^p) = a^(-p).
In you case 8=2^3, 1/8 = 1/(2^3) = 2^(-3)